Two electric charges 2 µC and –1 µC are placed at a distance of 20 cm from each


other in vacuum. Locate the point on the line joining these two charges outside the


region between them at which the electric potential is zero with reference to the


positive charge.

1
Expert's answer
2022-02-25T11:15:45-0500

"\varphi=0,"

"\varphi=\varphi_1+\varphi_2=\frac{kq_1}x+\frac{kq_2}{x-d}=0,"

"x=\frac d{1+\frac{q_2}{q_1}}=40~cm."